OpenStudy (anonymous):

There is a 72% chance that Katie will get married, a 46% chance that she will graduate college, and a 60% chance that she will marry and graduate. What is the chance that she will get married or graduate?

OpenStudy (amistre64):

theres a formula for this ... P(AuB) = P(A) + P(B) - P(AnB)

OpenStudy (muzzack):

what is the formula called @amistre64

OpenStudy (amistre64):

i call it fred .... not sure if it has a technical name

OpenStudy (muzzack):

ok well @amistre64 just gave you the formula now plug in the numbers

OpenStudy (muzzack):

what does P represent @amistre64

OpenStudy (amistre64):

 notice that in the venn diagram: A = 1+2 ; B = 2+3 ; and AnB = 2 A or B is usually thought of as the union of the set A with the set B. 1+2 + 2 + 3 is not an accurate representation of the union of the sets, its heavy by 2 (its counted twice) to adjust for this we have to subtract out the repetition